OpenStudy (anonymous):

f(x)=3x+2 Find the inverse function

OpenStudy (anonymous):

inverse functions means switch the x and the y: y = 3x + 2 x = 3y + 2 then solve for y x-2 = 3y (x-2)/3 = y